OpenStudy (anonymous):

reduce each fraction to its lowest terms. 21/28 65/70 18/42

OpenStudy (anonymous):

ok...can we start with the first one? 21/28

OpenStudy (anonymous):

21/28 divide both by 7 21/7 = 3 28/7 = 4 21/28 = 3/4

OpenStudy (anonymous):

what is the common number that both numbers are divisible by?

OpenStudy (anonymous):

go to google and type 21/28 = and it will give you the answer and if the answer to your fraction is lower then put them in order :)